OpenStudy (anonymous):

4x=-2y+26 solve for y

OpenStudy (anonymous):

\[ 4x = -2y + 26 \]You can cancel a 2 from each side and you get: \[ 2x = -y + 13 \rightarrow \boxed {y = 13 - 2x}. \]

OpenStudy (anonymous):

thank you~ i get it now :D

OpenStudy (anonymous):

No problem. We are here to help. I'm glad you understand now! :D (+thanks for the medal)

OpenStudy (fellowroot):